OpenStudy (qwertty123):

Does anyone know how to use a Casio fx-97550GII?

OpenStudy (qwertty123):

I need help using this Calc.

NvidiaIntely (nvidiaintely):

Check this out dude: https://www.youtube.com/watch?v=t6-HVcoQ5i4 yw

NvidiaIntely (nvidiaintely):

Let me know if you need more help with it (:

OpenStudy (qwertty123):

Okay.. lol I am not a dude! ^.^

NvidiaIntely (nvidiaintely):

oh okay.

OpenStudy (qwertty123):

Yup! ^.^ Me a Woman! xD Anyways let me see how this thing works..

OpenStudy (will.h):

That'sy calculator what do you need to know?

OpenStudy (will.h):

My*

OpenStudy (google):

\[2meny-ppl \frac{ viewing }{ this } \left(\begin{matrix}Question \\ ?\end{matrix}\right)\]

OpenStudy (qwertty123):

How to use it! I am so confused!

OpenStudy (mathmate):

If you have lost your instruction book, here it is: http://support.casio.com/manualfile.php?rgn=5&cid=004002013 Otherwise, try to post what you'd like to do but have difficulty with.

OpenStudy (qwertty123):

So I hit Menu how do you put an equation in it?

OpenStudy (will.h):

Hit menu and set the system in SAT

OpenStudy (qwertty123):

I get a whole bunch of different options...

NvidiaIntely (nvidiaintely):

did you watch this video? https://www.youtube.com/watch?v=t6-HVcoQ5i4

OpenStudy (qwertty123):

okay did that...

OpenStudy (will.h):

Sat is the original clear system now u can use it well

OpenStudy (will.h):

Other questions?

OpenStudy (qwertty123):

I pmed you..

OpenStudy (will.h):

Am on phone I hadn't notice.

OpenStudy (google):

wow... im sure there is an app

OpenStudy (google):

why would someone use an old calculator like that, theres gotta be and app

OpenStudy (will.h):

No the website works well

OpenStudy (will.h):

Oh thought u were talking to me

OpenStudy (qwertty123):

LOL it isn't old!

OpenStudy (google):

qwertty do you have a smartphone?

OpenStudy (qwertty123):

Why do you want to know?

OpenStudy (google):

if you have an android theres a free apk that lets you use the full version calculator on your phone

OpenStudy (google):

i have a link that downloads it

OpenStudy (google):

if you use that link on your phone it downloads the calculator

OpenStudy (qwertty123):

oh okay.... I got the Calc to work! ^.^

OpenStudy (google):

ok lol idk if you wanted to bring your pc everywhere

OpenStudy (qwertty123):

Thanks everyone...